how do you find the recursive formula of the sequence 1,1,2,3,5,8...?
Answer by jim_thompson5910(35256)

This is the fibonacci sequence. The first two terms are given as 1,1

The third term is then the sum of the previous two terms: 1+1 = 2

The fourth term is then the sum of the previous two terms: 1+2 = 3

The fifth term is then the sum of the previous two terms: 2+3 = 5

The sixth term is then the sum of the previous two terms: 3+5 = 8

and so on...

So the recursive formula is




--------------------------------------------------------------------------------------------------------------
If you need more help, email me at jim_thompson5910@hotmail.com

Also, please consider visiting my website: http://www.freewebs.com/jimthompson5910/home.html and making a donation. Thank you

Jim
--------------------------------------------------------------------------------------------------------------